OpenStudy (anonymous):

14√150/7√2 a:7√3 b.10√2 c.√150 d.10√3

OpenStudy (anonymous):

14√150/7√2 = 14√(25 * 6)/7√2 = 70√6/7√2 = 10√6/√2 = 10 *√3 * √2/√2 = 10√3

OpenStudy (anonymous):

10√3

OpenStudy (anonymous):

i don't understand how you got 14√(25 * 6) that part confuses me

OpenStudy (anonymous):

You see, You basically split the factors of 150. 25 * 4 = 150

OpenStudy (anonymous):

I mean 25 * 6 = 150

OpenStudy (anonymous):

oh! i get it now thank you so much!

OpenStudy (anonymous):

You're welcome :)